Nova Company’s total overhead cost at various levels of activity are presented below:

Month Machine-
Hours
Total
Overhead
Cost
April 52,000 $ 182,640
May 42,000 $ 158,940
June 62,000 $ 206,340
July 72,000 $ 230,040

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 42,000 machine-hour level of activity is:

Utilities (variable) $ 54,600
Supervisory salaries (fixed) 44,000
Maintenance (mixed) 60,340
Total overhead cost $ 158,940

Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements.

Required:

1. Estimate how much of the $230,040 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $230,040 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.)

2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX.

3. Express the company’s total overhead cost in the form Y = a + bX.

4. What total overhead cost would you expect to be incurred at an activity level of 47,000 machine-hours?
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Month Machine hours Total
Overhead
Cost
Utilities at 1.30 per unit Supervisory salaries Maintenance = Total - utilities - Sal
April               52,000.00            182,640.00              67,600.00           44,000.00           71,040.00
May               42,000.00            158,940.00              54,600.00           44,000.00           60,340.00
June               62,000.00            206,340.00              80,600.00           44,000.00           81,740.00
July               72,000.00            230,040.00              93,600.00           44,000.00           92,440.00
1) 92,440
2)
Cost at highest level of activity               92,440.00
Cost at Lowest level of activity               60,340.00
Highest level of activity               72,000.00
Lowest Level of activity               42,000.00
Variable cost per unit = (92440- 60340)/(72000 - 42000)
Variable cost per unit = 32,100/30,000
Variable cost per unit = $1.07
Fixed Costs = 60340- 42000*1.07
Fixed Costs = $15,400
Y = 15,400 + 1.07x
3) Y = 59,400 + 2.37x
Since fixed cost = 44000 + 15400
Variable cost = 1.30 + 1.07
4) total overhead cost = 59,400 + 47000 * 2.37 = $170,790